OpenStudy (anonymous):

find f(x+h) - f(x)/ h for the function f(x)=x^2 - 2x find f(x+h) - f(x)/ h for the function f(x)=x^2 - 2x @Mathematics

OpenStudy (anonymous):

please show steps

OpenStudy (across):

\[\lim_{h\to0}\frac{f(x+h)-f(x)}{h}\]of\[f(x)=x^2-2x\]Is this what you're asking?

OpenStudy (anonymous):

yes

OpenStudy (across):

All you have to do is plug the formula in. It's that easy:\[\lim_{h\to0}\frac{f(x+h)-f(x)}{h}=\lim_{h\to0}\frac{(x+h)^2-2(x+h)-x^2+2x}{h}=\]\[\lim_{h\to0}\frac{x^2+2xh+h^2-2x-2h-x^2+2x}{h}=\]\[\lim_{h\to0}\frac{2xh+h^2-2h}{h}=\lim_{h\to0}\frac{h(2x+h-2)}{h}=\lim_{h\to0}(2x+h-2)=2x-2\]

OpenStudy (across):

This is the definition of the derivative.